OpenStudy (anonymous):

how would i simplify (x-y)^2/y^2-x^2 using -1 as a factor

OpenStudy (anonymous):

(x-y)^2 = (x-y)(x-y) = -(y-x)(x-y) y^2-x^2 = (y-x)(y+x) so -(y-x)(x-y)/(y-x)(y+x) = -(x-y)/(y+x)